OpenStudy (anonymous):

The volume of a cone is 300 cubic kilometers. If the radius of the base is 7 kilometers, what is the height of the cone? Use for . Show your work and round to the nearest tenth

jhonyy9 (jhonyy9):

so do you know the formula for volum of a con ?

OpenStudy (anonymous):

yes its v=1/3*pi*r^2*h

OpenStudy (anonymous):

i also need help with another question sfter this one

jhonyy9 (jhonyy9):

so after you have wrote this formula for volum of a con can you calcule the height of con ?

OpenStudy (anonymous):

NO I DONT UNDERSTAND HOW TO DO THAT

jhonyy9 (jhonyy9):

ok from this formula of volum what is pi * r^2 ?

OpenStudy (anonymous):

YEP BUT IS THAT HOW TO FIND THE HEIGHT???????????????????

jhonyy9 (jhonyy9):

ok i know but these steps just for you can understanding easy it

OpenStudy (anonymous):

I KNOW HOW TO DO THE REGULAR FORMULA JUST NOW KNOWING HOW TO FIGURE OUT THE HEIGHT IS WAT IM SO CONFUSHED OF AND IT MIGHT LOOK LIKE IM MAD BUT IM NOT JUST FRUSTUACTED

jhonyy9 (jhonyy9):

ok so from this formula of volum you can calcule easy the height of con V=1/3 *(pi*r^2 *h) multiplie both sides by 3 than will get 3V =pi*r^2 *h divide both sides by pi*r^2 than will get 3V ------ = h pi*r^2 ok ?

jhonyy9 (jhonyy9):

can you calcule it now ?

OpenStudy (anonymous):

YEPPPP

jhonyy9 (jhonyy9):

ok good luck bye